OpenStudy (anonymous):

A drama club sold 120 child tickets and 80 adult tickets to a play. A child ticket cost $3 less than an adult ticket. The club collected a total of $640 from ticket sales. What is the cost of an adult ticket and a child ticket

OpenStudy (anonymous):

cost total=640 #child tickets=120 #adult tickets=80 c=child ticket cost a=adult ticket cost c+3=a 640=120c+80a subbing in a 640=120c+80(c+3) 640=120c+80c+240 200c=400 c=2 dollars 2+3=a a=5 dollars